One of the two numbers is greater than the other. if the sum of their reciprocals is 10/20, what are the number?
Found 2 solutions by josgarithmetic, Alan3354:
Answer by josgarithmetic(39630)
Answer by Alan3354(69443)

One of the two numbers is greater than the other. if the sum of their reciprocals is 10/20, what are the number?
---------------------------
1/x + 1/y = 10/20 = 1/2
(x+y)/xy = 1/2
2x + 2y = xy
===========
x = 3:
6 + 2y = 3y
y = 6
==================
x = 5:
10 + 2y = 10y
y = 10/8 = 5/4
===========
No unique solution.